# Proactive Agents for Multi-Turn Text-to-Image Generation Under Uncertainty

### v1

@@ Abstract
User prompts for generative AI models are often underspecified, leading to sub-optimal responses. This problem is particularly evident in text-to-image (T2I) generation, where users commonly struggle to articulate their precise intent. This disconnect between the user’s vision and the model’s interpretation often forces users to painstakingly and repeatedly refine their prompts. To address this, we propose a design for proactive T2I agents equipped with an interface to (1) actively ask clarification questions when uncertain, and (2) present their understanding of user intent as an understandable belief graph that a user can edit. We build simple prototypes for such agents and verify their effectiveness through both human studies and automated evaluation. We observed that at least 90% of human subjects found these agents and their belief graphs helpful for their T2I workflow. Moreover, we develop a scalable automated evaluation approach using two agents, one with a ground truth image and the other tries to ask as few questions as possible to align with the ground truth. On DesignBench, a benchmark we created for artists and designers, the COCO dataset (Lin et al., 2014), and ImageInWords (Garg et al., 2024), we observed that these T2I agents were able to ask informative questions and elicit crucial information to achieve successful alignment with at least 2 times higher VQAScore (Lin et al., 2024) than the standard single-turn T2I generation. Demo: https://github.com/google-deepmind/proactive_t2i_agents.
Figure 1: A proactive text-to-image agent interface that clarifies prompts, updates them based on user responses, and expresses its uncertainty and understanding as an editable belief graph.
@@ 1 Introduction
A fundamental challenge in the development of AI agents is how to foster effective and efficient multi-turn communication and collaboration with human users to achieve user-defined goals, especially when faced with the common issue of vague or incomplete instructions from humans. We focus specifically on text-to-image (T2I) generation, where recent advancements (Baldridge et al., 2024; Betker et al., 2023; Podell et al., 2023; Yu et al., 2023) have enabled the creation of stunning images from complex text descriptions. However, users often struggle to describe the image they would like to generate in a way that T2I systems can fully understand. This leads to unsatisfactory results and repeated iterations of prompts.
The prompt underspecification problem arises from the inherent ambiguity of natural language, the different assumptions that humans make and the vast space of potential images that can be generated from a single prompt (Hutchinson et al., 2022). Imagine a prompt generate an image of a rabbit next to a cat. This seemingly simple prompt leaves many important aspects underspecified: What kind of rabbit? What color is the cat? What is their relative positions? What is the background? While a T2I model can generate an image with a rabbit and a cat in it, it is unlikely that the image captures the specific details a specific user has in mind. For example, people in Holland might assume it is common for rabbits to have lop ears, but people in New England might expect to see a cottontail rabbit with straight ears. The combination of all these factors can lead to a frustrating cycle of trial-and-error, with the user repeatedly refining their prompt in an attempt to steer the model towards the desired output (Vodrahalli and Zou, 2024; Huang et al., 2024; Sun and Guo, 2023).
@@
Instead of relying on passive T2I models that simply generate images based on potentially vague user instructions, we pursue a quest for agency in T2I generation. The T2I agents should actively engage with human users to provide a collaborative and interactive experience for image creation. We envision that these T2I agents will be able to (1) express and visualize their beliefs and uncertainty about user intents, (2) allow human users to directly control their beliefs beyond just text descriptions, and (3) proactively seek clarification from the human user to iteratively align their understanding with what the human user intends to generate.
In this work, we develop simple prototypes of such agents. At the core of those agent prototypes, we build in a graph-based symbolic belief state, named belief graph, for agents to understand its own uncertainty about possible entities (e.g., rabbit) that might appear in the image, attributes of entities (e.g., rabbit’s color), relations between entities and so on. Given a user prompt, we use an LLM and constrain its generation to the graph structure of beliefs, which include probability estimates on the appearance of entities and the possible values for attributes and relations. Figure 1 illustrates the interface and features of the prototypes. In particular, the agent can ask questions based on its uncertainty. For example, a very simple strategy is to find the most uncertain attribute of an entity (e.g., rabbit’s color) and use an LLM to phrase a question about the attribute (e.g., What is the color of the rabbit?). The agent can also guide users to directly edit uncertain items in the graph.
@@
Based on user answers to agent questions or direct edits in the belief graph, the agent uses an LLM to update the prompt. It also transitions to a new belief graph by modifying the uncertainty of items clarified by the user. Based on the updated prompt, the agent calls an off-the-shelf T2I model to generate images. The structure of our agent prototypes is highly modular, making it easy to improve each component individually, e.g., changing the strategy for asking questions, updating the belief graph construction method, and switching to better LLMs or T2I models when they become available.
To evaluate the utility of our agent prototypes, we conduct both automatic evaluations and human studies. We develop automatic evaluation pipelines to assess the effectiveness and efficiency of the T2I agents when interacting with simulated users with underspecified prompts answering questions based on their pre-fixed intents. The human studies aim to understand how helpful simple T2I agents can be and evaluate how good the agents’ questions and generated images are.
@@
We also create a hand-curated benchmark called DesignBench which contains aesthetic scenes with multiple entities and interactions between entities; it also contains both a short and long caption. DesignBench features diversity between photo-realism, animation and multiple styles allowing a robust testing with the use case of artists and designers in mind.
@@
We run automatic evaluations on DesignBench, the COCO dataset (Lin et al., 2014) and ImageInWords (Garg et al., 2024). Results show that our agents can achieve at least 2 times higher VQAScore (Lin et al., 2024) than the traditional single-turn T2I generation within 5 turns of interaction. In our human study, over 90% human subjects expect proactive clarifications to be helpful, about 85% find belief graphs helpful, and 58% think the question asking feature of agents could deliver value to their work very soon, or immediately. Participants prefer images generated by our simple agents over those of single-turn T2I systems in more than 80% cases of the 550 prompt-image pairs used in the human study.
Our contributions: (1) the first explainable and controllable belief graph used for T2I, (2) novel design and prototypes for T2I agents that adaptively ask clarification questions and present belief graphs; (3) a new automatic evaluation pipeline with simulated users to assess question-asking skills of T2I agents; and (4) DesignBench: a new T2I agent benchmark. [^1] Code and DesignBench can be found at https://github.com/google-deepmind/proactive_t2i_agents. Appendix B details the novelty.
@@ 2 Related work
In machine learning and statistics, efficient data acquisition has been extensively studied for many problems, including active learning (Cohn et al., 1996; Settles, 2009; Houlsby et al., 2011; Gal et al., 2017; Ren et al., 2021; Wang et al., 2018), Bayesian optimization (Garnett, 2023; Kushner, 1964; Moc̆kus, 1974; Auer, 2002; Srinivas et al., 2010; Hennig and Schuler, 2012; Wang and Jegelka, 2017; Wang et al., 2024b), reinforcement learning (Kaelbling et al., 1996; Ghavamzadeh et al., 2015; Sutton, 2018) and experimental design (Chaloner and Verdinelli, 1995; Kirk, 2009). We reckon that T2I agents should also be capable of actively seeking important information from human users to quickly reduce uncertainty (Wang et al., 2024c) and generate satisfying images. In §E, we detail the implementation of action selection strategies for our T2I agents.
@@
In human-computer interaction, researchers have been extensively studying how to best enable Human-AI interaction especially from user experience perspectives (Norman, 1994; Höök, 2000; Amershi et al., 2019; Cai et al., 2019; Viégas and Wattenberg, 2023; Chen et al., 2024; Yang et al., 2020; Kim et al., 2023). Interface design for AI is becoming increasingly challenging due to the lack of transparency (Viégas and Wattenberg, 2023; Chen et al., 2024), uncertainty about AI capability and complex outputs (Yang et al., 2020). We aim to build user-friendly agents, and an indispensable component is their interface to enable them to effectively act and observe, as detailed in §G.
@@
Interpretebaility. Surfacing an agent’s belief overlaps with interpretability as both aim to understand model or agent’s internal. Some methods leverage LLM’s natural language interface to surface their reasoning (e.g., chain of thought (Wei et al., 2023a)), sometime interactively (Wang et al., 2024a). While these approaches make accessible explanations, whether the explanations represent truth has been questioned (Lanham et al., 2023; Wei et al., 2023b; Chen et al., 2023). Some studies indicate explanations generated by the LLMs may not entail the models’ predictions nor be factually grounded in the input, even on simple tasks with extractive explanations (Ye and Durrett, 2022). In this work, the belief graph does not correspond to the distribution over outputs of the T2I model itself conditioned on the underspecified prompt. Instead, the belief graph is designed to align with the distribution over images generated by the agent, since the agent can construct detailed prompts according to its belief, and feed them into a high-quality T2I model.
@@
Text-to-Image (T2I) generation. Text-to-image prompts can be ambiguous, subjective (Hutchinson et al., 2022), or challenging to represent visually (Wiles et al., 2024). Different users often have distinct requirements for image generation, including personal preferences (Wei et al., 2024), style constraints (Wang et al., 2023), and individual interpretations (Yin et al., 2019). To create images that better align with users’ specific needs and interpretations, it is essential to actively communicate and interact with the user to understand the user’s intent.
[…]
A core challenge in multi-turn T2I is consistency (Cheng et al., 2024a, b; Zeqiang et al., 2023). Hu et al. (2024) introduce Instruct-Imagen, which is a model that follows complex multi-modal instructions. AudioStudio (Cheng et al., 2024a) is a multi-turn T2I framework aimed at subject consistencies while generating diverse and coherent images. These consistency improvement methods can potentially be integrated into our T2I agents since they are highly modular. However, as an ablation, we only focus on the sequential decision making capability of agents to elicit user intents.
@@
Evaluating image-prompt alignment is important for T2I models. Relevant metrics can be embedding-based, such as CLIPScore (Hessel et al., 2022), ALIGNScore (Zha et al., 2023), VQA-based such as TIFA (Hu et al., 2023), DSG (Cho et al., 2023) and VQAScore (Lin et al., 2024), and captioning-based like LLMScore (Lu et al., 2023). Approaches such as PickScore (Kirstain et al., 2023), ImageReward (Xu et al., 2023) and HPS-v2 (Wu et al., 2023) finetune models on human ratings to devise a metric that aligns with human preferences. Recently, diversity of generated images (Naeem et al., 2020) is also becoming an important metric of measurement to track progress, especially in the geo-cultural context (Kannen et al., 2024; Hall et al., 2024).
In this work, we develop an automatic approach to evaluate agent-user conversations. We adopt VQAScore (Lin et al., 2024) to evaluate the alignment between a ground truth prompt and an image generated by an agent after interactions with a simulated user. Other T2I metrics can also be used.
@@
Prompt expansion is a widely known technique to improve image generation (Betker et al., 2023). ImageinWords (Garg et al., 2024) proposes to obtain high-quality hyper-detailed captions for images, which significantly improve quality of image generation. Datta et al. (2024) present a generic prompt expansion framework used along Text-to-Image generation and show an increase in user satisfaction through human study. While our work can be viewed as a method to adaptively expand a T2I prompt based on user feedback [^2] Samples from the agent belief can be used to construct expanded prompts., evaluating our method as a prompt expansion tool is outside of our scope.
@@ 3 Background
The belief graph in our work is closely related to symbolic world representations.
World states.
In classical AI, researchers use symbolic representations to describe the world state (McCarthy and Hayes, 1969; Minsky, 1974, 1988; Pasula et al., 2007; Kaelbling and Lozano-Pérez, 2011). For example, in the blocks world (Ginsberg and Smith, 1988; Gupta and Nau, 1992; Alkhazraji et al., 2020), a state can be
$is\_block(a)\wedge is\_red(a)\wedge on\_table(a)\wedge is\_block(b)\wedge is\_blue(b)\wedge on(b,a),$
describing that there are a red block and a blue block, referred to as $a$ and $b$, block $a$ is on a table, and block $b$ is on $a$. Such world states must include entities (e.g., $a$ and $b$), their attributes (e.g., position $on\_table$, characteristics $is\_block$) and relations (e.g., $on(b,a)$) which are critical for enabling a robot to know and act in the world.
In linguistics, Davidson (1965, 1967b, 1967a) introduce logic-based formalisms of meanings of sentences. The semantics of a sentence is decomposed to a set of atomic propositions, such that no propositions can be added or removed from the set to represent the meaning of the sentence. (Cho et al., 2023) propose Davidsonian Scene Graph (DSG) which represent an image description as a set of atomic propositions (and corresponding questions about each proposition) to evaluate T2I alignment.
@@
Belief states.
Term “belief state” (Nilsson, 1986; Kaelbling et al., 1998) has been used to describe a distribution over states. E.g., for block $a$, we might have $p(on\_table(a))=0.5$ and $p(\neg on\_table(a))=0.5$, which means the agent is unsure whether the block is on a table. To represent the T2I agent’s belief on which image to generate, we need to consider the distribution over all possible “worlds” in which the picture can be taken. This distribution can be described by the probabilities that an entity appears in the picture, an attribute gets assigned a certain value, etc.
A core difference between our belief graphs and classic belief states is that belief graphs do not need pre-defined predicates (e.g., $on(b,a)$), but instead automatically produce useful predicates using LLMs. This makes belief graphs much more generalizable across domains. More details in §B.

### v2

@@ Abstract
User prompts for generative AI models are often underspecified, leading to a misalignment between the user intent and models’ understanding. As a result, users commonly have to painstakingly refine their prompts. We study this alignment problem in text-to-image (T2I) generation and propose a prototype for proactive T2I agents equipped with an interface to (1) actively ask clarification questions when uncertain, and (2) present their uncertainty about user intent as an understandable and editable belief graph. We build simple prototypes for such agents and propose a new scalable and automated evaluation approach using two agents, one with a ground truth intent (an image) while the other tries to ask as few questions as possible to align with the ground truth. We experiment over three image-text datasets: ImageInWords (Garg et al., 2024), COCO (Lin et al., 2014) and DesignBench, a benchmark we curated with strong artistic and design elements. Experiments over the three datasets demonstrate the proposed T2I agents’ ability to ask informative questions and elicit crucial information to achieve successful alignment with at least 2 times higher VQAScore (Lin et al., 2024) than the standard T2I generation. Moreover, we conducted human studies and observed that at least 90% of human subjects found these agents and their belief graphs helpful for their T2I workflow, highlighting the effectiveness of our approach. Code and DesignBench can be found at https://github.com/google-deepmind/proactive_t2i_agents.
Figure 1: Our proactive T2I agent clarifies a user prompt with questions, incorporates user feedback, and expresses its uncertainty and understanding as an editable belief graph.
@@ 1 Introduction
A fundamental challenge in the development of AI agents is how to foster effective and efficient multi-turn communication and collaboration with human users to achieve user-defined goals, especially when faced with the common issue of vague or incomplete instructions from humans.
As a testbed for the communication problem, this work uses text-to-image (T2I) generation, where recent advancements (Baldridge et al., 2024; Betker et al., 2023; Podell et al., 2023; Yu et al., 2023) have enabled the creation of stunning images from complex text descriptions. T2I exemplifies the challenge of moving from low-information (prompts) to high-information (images), forcing models to make implicit assumptions to resolve prompt uncertainty. Moreover, T2I offers well-defined metrics and a clear problem scope, enabling comprehensive automatic evaluation.
Interpretations of prompts can vary significantly across diverse populations, groups, and cultures, resulting in misalignment between user intents and model outputs. For example, the prompt “A breakfast plate” can evoke a wide range of mental images depending on cultural background. While a typical T2I model might generate a generic breakfast plate, it is unlikely to produce a culturally appropriate cuisine without further information about the user’s background and desired cuisine. These model assumptions can lead to user frustration and dissatisfaction, repeated prompt adjustments, while also potentially propagating bias (Kannen et al., 2024; Basu et al., 2023) and causing offense (Bianchi et al., 2023).
Beyond cultural contexts and differences, similar challenges arise with complex scenes. Consider the prompt “Image of a rabbit near a cat.” Standard T2I models will likely generate an image containing both animals, however, it is unlikely the generated images will capture the specific details and spatial relationships the user envisions. These combined factors often result in a frustrating cycle of trial and error, as users repeatedly refine their prompts in an attempt to guide the model toward the desired output (Vodrahalli and Zou, 2024; Huang et al., 2024; Sun and Guo, 2023).
@@
In this paper we pursue a quest for agency in T2I generation with the aim to mitigate the issues that arise from passive T2I models. In the proposed setup, T2I agents actively engage with human users to provide a collaborative and interactive experience for image creation. We envision that these T2I agents will be able to (1) express and visualize their beliefs and uncertainty about user intents, (2) allow human users to directly control agents’ beliefs beyond just text descriptions, and (3) proactively seek clarification from the human user to iteratively align their understanding with what the human user intends to generate.
At the core of these agents, we build in a graph-based symbolic belief state, termed the belief graph, that allows agents to express uncertainty about entities (e.g., “breakfast”, “table”) that may appear in the image, attributes of entities (e.g., cuisine type) and relations between entities (e.g., “breakfast” being on “table”). When a user provides a prompt, we use an LLM and constrain its generation to adhere to the belief graph data structure, which includes probability estimates for entity presence and potential values for attributes and relations.
For example, for a prompt related to “homemade breakfast on a wooden table” in Figure 1, the agent’s belief graph might indicate a 9% probability that the cuisine is American, while also recognizing other possibilities like Indian or Faroese. The belief graph also anticipates entities not explicitly mentioned but the user might want. Even though the user only specifies “homemade breakfast”, the agent might be 35% confident that a fork will be in the generated images. The expression of uncertainty enables the agent to inspire new ideas and to guide users in editing uncertain items within the graph.
However, uncertainty alone is an insufficient trigger for clarification, as not all uncertain elements are equally consequential. For instance, the breakfast’s cuisine type is likely more impactful on the generated images than the table’s wood type. To address this, we embed importance scores alongside probabilities within the belief graph for each entity, attribute, and relation.
Based on both uncertainty and importance, the agent proactively asks clarification questions. It can, for instance, identify the most uncertain attribute of an important entity (e.g., cuisine type of “breakfast”). An LLM is then used to formulate a relevant question (e.g., “Which type of cuisine should be pictured?”). To make answering easier, the agent provides likely options for users to select from (e.g., American, Indian, European, Chinese), though users can also type their own answers directly.
@@
Conditioned on user answers and edits in the belief graph, the agent uses an LLM to update the prompt and transitions to a new belief graph by modifying the uncertainty of items clarified by the user. With the updated prompt, the agent calls an off-the-shelf T2I model to generate images.
The structure of our agent prototypes is highly modular, making it easy to improve each component individually, e.g., changing the strategy for asking questions, updating the belief graph construction method, and switching to better LLMs or T2I models when they become available.
To address the lack of evaluation standards for proactive T2I agents, we develop a scalable and automatic evaluation pipeline with a collection of metrics. The system first constructs a simulated user whose ground truth intent is a prefixed image, and translates the image to a short, underspecified text description. We then let a proactive agent and the simulated user play a multi-turn question answering game. The proactive agent aims to collect as much information as possible to generate images close to the ground truth, while the simulated user provides concise answers to agent questions. We then use text and image similarity metrics to evaluate the generated image against the goal image. We observe a strong correlation between the automatic evaluation results and human evaluation results, highlighting our pipeline’s reliability in assessing T2I agents.
@@
To better represent intents of diverse artists and designers (a key group of T2I users), we have created DesignBench, a carefully hand-curated benchmark. This benchmark contains aesthetic scenes with multiple entities and interactions between them. It also provides both short and long captions that standardize simulated users in our evaluation pipeline. DesignBench features diversity between photo-realism, animation and multiple styles, allowing a robust testing with the use case of artists and designers in mind.
@@
We run both automatic and human evaluations on DesignBench, the COCO dataset (Lin et al., 2014) and ImageInWords (Garg et al., 2024). Our proposed agents achieve at least 2 times higher VQAScore (Lin et al., 2024) than traditional single-turn T2I generation within just 5 turns of interaction. Additionally we conduct a human user study in which over 90% human subjects expect proactive clarifications to be helpful, about 85% find belief graphs helpful, and 58% think the question asking feature of agents could deliver value to their work very soon, or immediately. Participants prefer images generated by our agents over those of single-turn T2I systems in more than 80% cases of the 550 prompt-image pairs used in the human study. This shows that our simple framework, which uses an off the shelf model without fine-tuning, can provide significant gains.
We believe our approach, where agents actively gather information from the user to reduce uncertainty, will empower a diverse range of users to leverage generative models for their needs and make these models safer and more responsible. By enabling direct control via interaction and understandable belief graphs, we move beyond prompt crafting and towards a future where tools like T2I are more accessible to people without special knowledge of writing prompts.
@@ 2 Related Work
In machine learning and statistics, efficient data acquisition has been extensively studied for many problems, including active learning (Cohn et al., 1996; Settles, 2009; Houlsby et al., 2011; Gal et al., 2017; Ren et al., 2021; Wang et al., 2018), Bayesian optimization (Garnett, 2023; Kushner, 1964; Moc̆kus, 1974; Auer, 2002; Srinivas et al., 2010; Hennig and Schuler, 2012; Wang and Jegelka, 2017; Wang et al., 2024b), reinforcement learning (Kaelbling et al., 1996; Ghavamzadeh et al., 2015; Sutton, 2018) and experimental design (Chaloner and Verdinelli, 1995; Kirk, 2009). We reckon that T2I agents should also be capable of actively seeking important information from human users to quickly reduce uncertainty (Wang et al., 2024c) and generate satisfying images. In §D, we detail the implementation of action selection strategies for our T2I agents.
@@
In human-computer interaction, researchers have been extensively studying how to best enable Human-AI interaction especially from user experience perspectives (Norman, 1994; Höök, 2000; Amershi et al., 2019; Cai et al., 2019; Viégas and Wattenberg, 2023; Chen et al., 2024; Yang et al., 2020; Kim et al., 2023). Interface design for AI is becoming increasingly challenging due to the lack of transparency (Viégas and Wattenberg, 2023; Chen et al., 2024), uncertainty about AI capability and complex outputs (Yang et al., 2020). We aim to build user-friendly agents, and an indispensable component is their interface to enable them to effectively act and observe, as detailed in §F.
@@
Interpretability. Surfacing an agent’s belief overlaps with interpretability as both aim to understand model or agent’s internal state. Some methods leverage LLM’s natural language interface to surface their reasoning (e.g., chain of thought (Wei et al., 2023a)), sometime interactively (Wang et al., 2024a). While these approaches make accessible explanations, whether the explanations represent the truth has been questioned (Lanham et al., 2023; Wei et al., 2023b; Chen et al., 2023). Some studies indicate explanations generated by the LLMs may not entail the models’ predictions nor be factually grounded in the input, even on simple tasks with extractive explanations (Ye and Durrett, 2022). In this work, the belief graph does not correspond to the distribution over outputs of the T2I model itself conditioned on the underspecified prompt. Instead, the belief graph is designed to align with the distribution over image-prompts generated by the agent, since the agent can construct detailed prompts according to its belief, and feed them into a high-quality T2I model.
@@
Text-to-Image (T2I) generation. Text-to-image prompts can be ambiguous, subjective (Hutchinson et al., 2022), or challenging to represent visually (Wiles et al., 2024). Different users often have distinct requirements for image generation, including personal preferences (Wei et al., 2024), style constraints (Wang et al., 2023), and individual interpretations (Yin et al., 2019). To create images that better align with users’ specific needs and interpretations, it is essential to actively communicate and interact with the user to understand the user’s intent. Other work (Mehrabi et al., 2022) proposes to use clarification questions to resolve ambiguities (a prompt has multiple meanings), but our aim is to use clarification questions to resolve under-specification (the prompt is not ambiguous, but lacks information to fully describe the image).
@@
T2I models are evaluated by analyzing image-prompt alignment using metrics that can be embedding-based, such as CLIPScore (Hessel et al., 2022), ALIGNScore (Zha et al., 2023), VQA-based such as TIFA (Hu et al., 2023), DSG (Cho et al., 2023) and VQAScore (Lin et al., 2024), and captioning-based like LLMScore (Lu et al., 2023). Approaches such as PickScore (Kirstain et al., 2023), ImageReward (Xu et al., 2023) and HPS-v2 (Wu et al., 2023) finetune models on human ratings to devise a metric that aligns with human preferences. Diversity of generated images (Naeem et al., 2020) is also used as a metric to track progress, especially in the geo-cultural context (Kannen et al., 2024; Hall et al., 2024). To evaluate our agent-user conversations and corresponding generated images, we develop an automatic approach using a variety of the metrics mentioned as well as collect human ratings on the generated images and generated conversations.
@@
Prompt expansion is a widely known technique to improve image generation (Betker et al., 2023). ImageinWords (Garg et al., 2024) proposes to obtain high-quality hyper-detailed captions for images, which significantly improve the quality of image generation. Datta et al. (2024) present a generic prompt expansion framework used alongside Text-to-Image generation and show an increase in user satisfaction through human study. While our work can be viewed as a method to adaptively expand a T2I prompt based on user feedback [^1] Samples from the agent belief can be used to construct expanded prompts., evaluating our method as a prompt expansion tool is outside of our scope.
@@ 3 Background
The belief graph in our work is closely related to symbolic world representations. More extensive related works on human-computer interaction, interpretability, prompt expansion are in §2.
World states.  Classical AI represents the world with symbols (McCarthy and Hayes, 1969; Minsky, 1974, 1988; Pasula et al., 2007; Kaelbling and Lozano-Pérez, 2011). For example, in the blocks world (Ginsberg and Smith, 1988; Gupta and Nau, 1992; Alkhazraji et al., 2020), a state can be $is\_block(a)\wedge is\_red(a)\wedge on\_table(a)\wedge is\_block(b)\wedge is\_blue(b)\wedge on(b,a),$ describing that there are a red block and a blue block, referred to as $a$ and $b$, block $a$ is on a table, and block $b$ is on $a$. Such world states include entities (e.g., $a$ and $b$), their attributes (e.g., position $on\_table$, characteristics $is\_block$) and relations (e.g., $on(b,a)$) which are critical for enabling a robot to know and act in the world. The predicates like $on\_table$ are pre-defined and hardcoded into robot systems.
In linguistics, Davidson (1965, 1967b, 1967a) introduce logic-based formalisms of meanings of sentences. The semantics of a sentence is decomposed to a set of atomic propositions, such that no propositions can be added or removed from the set to represent the meaning of the sentence. Cho et al. (2023) propose Davidsonian Scene Graph (DSG) which represents an image description as a set of atomic propositions (and corresponding questions about each proposition) to evaluate T2I alignment.
@@
Belief states.  Term “belief state” (Nilsson, 1986; Kaelbling et al., 1998) has been used to describe a distribution over states. E.g., for block $a$, we might have $p(on\_table(a))=0.5$ and $p(\neg on\_table(a))=0.5$, which means the agent is unsure whether the block is on a table. To represent the T2I agent’s belief on which image to generate, we need to consider the distribution over all possible “worlds” in which the picture can be taken. This distribution can be described by the probabilities that an entity appears in the picture, an attribute gets assigned a certain value, etc.
In contrast to classical belief states, our belief graphs incorporate importance scores tailored to generative AI tasks and leverage LLMs to automatically derive predicates from user descriptions. This approach makes the belief graphs highly adaptable to new tasks, eliminating the need for the rigid, hardcoded rules in traditional AI systems.
4 Belief Graphs
Instead of using hardcoded symbols in classic belief representations (Fikes and Nilsson, 1971) described in §3, We adopt a simple approach: leveraging LLMs as a tool to adaptively generate belief graphs from image descriptions of any length. While more sophisticated approaches, such as those using activations, can be explored as future work, this simple method allows for flexible belief graph generation across many types of images.
In a belief graph, nodes are entities and edges are relations between entities. Each entity is composed of its name, description, type, probability, importance score and attributes.
We have 3 types of entities: (a) explicit: entities mentioned in the prompt, (b) implicit: entities not mentioned in the prompt but likely to appear, e.g., fork in fig. 1, and (c) background entities, such as image style, time of day, location, which play important roles in constructing the image.
Each entity’s probability reflects its likelihood of appearing in the generated image, given the user’s prompt. This can be thought of as the output of a classifier. Explicit and background entities typically have probabilities close to 1 or 0, clearly indicating their presence or absence. Implicit entities, however, usually have probabilities between 0 and 1, reflecting their uncertain inclusion.
An entity’s importance score (normalized between 0 and 1) reflects its relevance and impact on the image. For example, in the “breakfast spread out across a wooden table” prompt in fig. 1, we care more about the specific breakfast items than the type of wood the table is made of. Therefore, “breakfast” should have a higher importance score than “table”. It is worth noting that different people have different judgments on the relative importance of elements in an image. In this work, we make a simplifying assumption that the agent doesn’t consider the distribution over people’s perceptions of importance scores.
An entity has multiple attributes, each with a name, a list of value-probability pairs (see “cuisine” in fig. 1) and an importance score. An attribute’s importance score shows its relevance to the image. For example, the cuisine of the breakfast might be more important than its exact size.
Relations are a special case of attribute that connects two entities, along with other information that an attribute can have. For each relation, the values in the list of value-probability pairs can be “part of”, “under”, “overlap”, etc., describing the spatial relations between the two entities. Besides explicit relations (e.g., breakfast on table), belief graph also includes potential implicit relations. For instance, the spatial relation between “fork” and “table” might have a name like “in the center of”.